OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) A certain bakery has found that the daily demand for bran muffins is 7700/p, where p is the price of a muffin in cents. The daily supply is 3p- 100. Find the price at which supply and demand are equal.

OpenStudy (anonymous):

guess you have to solve \[\frac{7700}{p}=3p-100\]

OpenStudy (anonymous):

google It

OpenStudy (anonymous):

start by multplying both sides by \(p\) and solve \[7700=3p^2-100p\]

OpenStudy (anonymous):

then yeah , you can google it http://www.wolframalpha.com/input/?i=7700%3D3p^2-100p

OpenStudy (anonymous):

by some miracle \[3p^2-100p-7700=0\] actually factors as \[(3p+110)(p-70)=0\]

OpenStudy (anonymous):

P=70 I'm 100% sure

OpenStudy (anonymous):

P=-100/3